Boots Roofing just paid its annual dividend of $1.85 a share. The firm recently announced that all future dividends will be increased by 3.5 percent annually. What is one share of this stock worth to you today if you require a 7 percent rate of return?
Answer #1

Div Po r-g Po= Price of Stock Div = Estimat ed Dividends for Next Period rRequired Rate of Return g Growth Rate

Div1 = $1.85 * (1 + 3.5%) = $1.91475

1.91475 Ро 0.07-0,035Answer Po=54.71